Find the distribution law oi te a 6. A software package consists of 11 programs, 7 of which must be upgraded. 5 programs are neednt be upgraded among chosen. randomy chosen for testing. Construct the distribution law for the number of programs that bilitu donsity function:
Answer #1

Ans:

Probability that a program need to be upgraded=7/11

Probability that a program need not to be upgraded=4/11

Use Binomial distribution with n=5 and p=4/11

P(x=k)=5Ck*(4/11)k*(7/11)5-k

where, k=0,1,2,3,4,5